Activity has been at a low level following the explosive events of early August. Hybrid earthquakes swarms continued during the early part of the reporting period. Rockfall activity replaced the hybrid events as the dominant form of seismic signal on August 20 until the end of the reporting period. Pyroclastic flows have continued to occur in the Gage's Valley with small flows also in the Tar River Valley and Mosquito Ghaut.
Activity declined from a very high level following the explosive events of early-August. The first few days of the reporting period showed a progressive decline in the level of pyroclastic flow activity, from moderate flows down Fort Ghaut reaching the end of Lovers Lane on 16 August to rockfall activity in Upper Gages on 21 August. Views of the dome on 18 and 19 August showed that growth is still concentrated above Gages in an area about 150m wide, comprising several massive blocks and spines. Theodolite measurements showed that one spine is at a height of 970m asl. Pyroclastic flow activity continued at a low level after 21 August. Most of the flows occurred in Gages Valley, although small flows occurred in the Tar River Valley on 29 August and intermittently in Mosquito Ghaut.
For the first four days of this period the activity continued to be characterised by intense hybrid swarms occurring at approximately 8 hour intervals. Then, on the 20th, hybrid activity declined and rockfall activity became dominant. Rockfalls were concentrated in "bursts" of activity roughly every 12 hours, although this activity was less regular than the hybrid swarms had been. It was noted that during the periods of more intense rockfall activity the background level on the RSAM plots rose, although nothing was observed on the drums. Detailed examination of the broadband record revealed tiny hybrid type signals among the noise between rockfalls. These were too small to trigger the networks but contributed to the background amplitude level. These mini-hybrids could correspond to magma entering the dome and causing the rockfall activity by making the dome swell. Towards the end of this period the number of long period rockfalls and VT's was increasing.
Table 1: Earthquake types
These earthquake counts are of events that triggered the broadband network's event recording system between 0000 and 2400 each day (local time).
Date VT Hybrid LP Dome RF LPRF* HYRF* 16 Aug 97 0 405 0 17 0 0 17 Aug 97 0 152 0 24 0 0 18 Aug 97 0 330 5 25 0 0 19 Aug 97 0 259 0 33 0 0 20 Aug 97 1 69 2 60 1 0 21 Aug 97 1 28 6 90 1 0 22 Aug 97 0 120 0 35 0 0 23 Aug 97 0 24 4 58 0 0 24 Aug 97 0 11 3 37 0 0 25 Aug 97 3 24 4 69 3 0 26 Aug 97 5 7 13 93 5 0 27 Aug 97 5 5 5 47 5 0 28 Aug 97 6 3 30 62 6 0 29 Aug 97 12 2 33 109 15 1 30 Aug 97 6 8 23 88 18 1
*LPRF: LP earthquake followed by rockfall signal. HYRF: Hybrid earthquake followed by rockfall signal. The LPs, hybrids and rockfalls in these signals are also counted in their respective columns.
A GPS occupation of BIGNET was made on 27 August (stations at Harris, Whites, Long Ground and O'Garra's). The occupation indicated continued northward motion of the sites at Whites and Long Ground. The baseline Harris-Whites has now shortened by over 4cm from its long-term mean length on which we had a 2-sigma confidence of 9mm. Shortening on the line began around the middle of March and has gradually accelerated. The last four measurements have all been shorter than 2-sigma below the mean. Long Ground has shown a broadly similar trend - its movement through 1996 to the middle of April 1997 was 3cm to the east. Since that time the movement direction has changed to the north-east and the site has moved by another 4cm. The O'Garra's pin had not been occupied since early-March 1997 due to the danger at the mouth of the White River. This occupation indicated around 3-4cm northward movement of the pin, although further occupations are necessary incase this was a single unreliable result. Lines between Whites, Long Ground and O'Garra's are unchanged. The most recent occupations of lines between Harris and Windy Hill, Dagenham and Brodericks also all appear to be stable.
EDM occupations were made on the Water Works - Lee's Yard, MVO - Lee's Yard and MVO - Garibaldi Hill lines. Slant distances on these lines remained generally stable throughout the reporting period. There is some slight shortening of the slant distance between Water Works and Lee's Yard. The erratic behaviour of the WW-LY line observed in the past appears to have been isolated to the period July 25 - August 3 and has not been observed since then. It is still unclear as to what the cause of these jumps were but they are generally not believed to be the result of elastic ground deformation.
Dome Volume Measurements
The programme of volume measurements continued with GPS-laser binocular surveys by helicopter of the pyroclastic flow deposits in Tuitt's Ghaut, the Farm/Trants area and the Fort Ghaut/Plymouth deposits. For completion the full series of deposit volume measurements are reported.
Location Million cubic meters Tar River Valley 13.55 Tar River Fan 15.72 Mosquito Ghaut 8.88 Tuitts Ghaut 1.68 Trant's/Farm 4.24 Fort Ghaut 29.83
The Trant's/Farm deposit is taken to be material lying to the north and east of the Tuitts Ghaut - Mosquito Ghaut confluence emplaced almost entirely during the 25 June collapse. Following 25 June the only contribution of deposit to this area was from pumice flows during the early-August explosive events. These formed thin lobes less than 2m thick and are at the limit of detection of the technique. The deposits are at their thickest along the channel of Pea Ghaut where they are over 20m and thin to around 5m over the bulk of the fan.
Debris accumulated very rapidly in Fort Ghaut during late-July and early-August. Volumetrically at least 90% of the material is contained in the upper reaches of Gages Valley and in the fan in the Glen Mhor area. The fan covers an area of around 0.25km2, is over 30m thick and has totally buried some of the houses at Glen Mhor. The margins of the ghaut are also buried apart from a small stretch just south of the Lower Soufriere. Flows down Fort Ghaut are now almost unconstrained once they emerge from the gap between Gages Mountain and Chances Peak.
A bathymetric survey of the sea bed around the Tar River Fan was carried out on 25 August using the MVO GPS equipment and the echo sounder on the police launch. No substantial changes were recorded since the previous survey on 21 March.
Photographs of the dome were taken from sites around the volcano during a clear period on 27 August to measure its growth rate.
Dust Trak sampling has been carried out at several sites around the island showing low values in the northern area and higher values in the Salem area. Values have slightly decreased compare to the previous week and they have been relatively stable during these two weeks. Results are shown in Table 2.
Table 2. PM10 (concentration in mg/m3)
Loc. 17/8/97 18/8/97 19/8/97 20/8/97 21/8/97 22/8/97 23/8/97 24/8/97 25/8/97 26/8/97 27/8/97 28/8/97 29/8/97 30/8/97 BC - - - - - - - - 0.034 - - - - - CB - - - - - - - - 0.03 P 0.04 - - - - CCSo - - - - 0.059 - - - - - - - - - Frith - 0.148 - - - - - - - - - - - - Fogarthy- - - - - - - - - - - - 0.005 - GPhel. - - 0.046 - - - - - - - - - - - Gpi - - - - - - - - P 0.077 - - - - - Gpo - - - - - - - - 0.04 - - - - - Lawyer's- - - - - - - - - - - - 0.007 - Mc - - - - - - - - - - 0.035 - - - MDF - - - - - - - - - - - - 0.026 - MVOi c 0.054 d 0.091 - - - - - - - - - - - - MVOo - - d 0.035 f 0.024 b 0.027 - d 0.039 f 0.050 - c 0.043 d 0.043 d 0.041 d 0.024 f 0.012 MVOpd - - a 0.07F - - - - - - - - - - - RBC - - - - 0.059 - - - - - 0.101 - - - RC - - 0.017 - - - - - 0.037 P 0.053 - - - - RM - - - - 0.042 - - - - - - - - - S.Pet. - - - - 0.029 - - - - - - - - - SCP - 0.12 0.056 - 0.08 0.099 0.093 - 1.067 0.080 0.125 - - - Sso - - 0.048 - - - 0.037 - 0.105 0.064 - - - - SSSi - 0.079 - - - - 0.014 - - - - - - - SSSo - 0.042 - - - - - - 0.064 P 0.169 0.048 - - - WB - - - - - - - - - - 0.033 - - -
These values represent the average reading over the sample duration. The mean duration of sampling is usually between 15 to 30 minutes. Samples taken over longer duration are labelled as follows:
- No samples taken i sample taken inside o sample taken outside F during fallout P personal sample (during activity) Key to locations: BC: Brades Church (North) CB: Carr's Bay (North) CCS: Catholic Church Shelter (Salem) GP: Gerald's Park Shelters (North) GPhel.: Gerald's Park heliport Mc: McChesney (Olveston) MDF: Montserrat Defense Force (Gerald's Bottom) MVO: Montserrat Volcano Observatory (Old Towne) MVOpd: MVO pool deck RBC: Royal bank of Canada (Olveston) RC: Red Cross (Woodland) RM: Radio Montserrat (Olveston) S.Pet.: S. Peters (North) SCP: Salem Car Park SS: Supermarket in Salem (RAMS EMDEE) SSS: Secondary School Shelter (Salem) WB: Woodland beach.
Monitoring of the average background levels of SO2 gas in inhabited areas was again initiated during this period. Between 14th August and 29th August SO2 diffusion tubes were left at the old MVO in Old Towne, at Lawyers water tower, at Fogarty and at the Montserrat Defence Force base. Over the two week period none of these sites was exposed to measurable levels of SO2.
The results are shown in Table 3.
Table 3. Sulphur dioxide diffusion tube results for the period between
14 and 29 August. Concentrations are in ppb.
Location SO2/ppb Lawyers water tower <0.02 Fogarthy <0.02 MDF base <0.02 MVO south <0.02
Rain water samples were collected at 3 locations around the island on 22 and 29 August, and results are shown in Table 3. These results show values at an acceptable level for the samples taken on the 22 but more acid values for the samples taken on the 29 with also a high concentrations of certain anions during this week.
Table 4. Rain water geochemistry, 22 Aug and 29 Aug 97.
Units: conductivity (mS/cm), total dissolved solids (g/l), fluoride (mg/l), chloride (mg/l), sulphate (mg/l), ND: not detected
Location Date Ph Cond. TDS Slphts Chlrds Flrds Lawyer's 22/8 6.68 0.054 0.027 ND 5.8 0.20 Fogarthy 22/8 6.58 0.065 0.032 3 8.0 0.25 MVO south 22/8 6.29 0.056 0.027 3 9.0 0.25 Lawyer's 29/8 3.09 0.417 0.209 3 52 0.95 Fogarthy 29/8 6.15 0.044 0.022 ND 7 0.10 MVO south 29/8 2.18 3.050 1.520 2 260 1.50
There was no ash to collect.
MVO Staff Changes
Rod Stewart (AWE Blacknest/BGS)
Dr Simon Young (BGS)
Richie Robertson (SRU, Trinidad)